From GlueXWiki
Jump to: navigation, search

August 25, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Engineering (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at EEL126 (cell#2) (Beni)
    • Results with new gas mixture: [6]
  5. Other: how to assemble a package


Participants: Eugene, Bill, Dave, Beni, Chris, Simon, and Lubomir.


- Dave: yesterday finished with stringing and taping wire frame #6, today did the position measurements and by the end of the day the wires will be glued. The cutter is now again operational and the techs will continue with cutting the foils for the cathode set #7; working also on the flaps for the cathode type 2 from set #6.

- HV cap problem (see pictures linked above). Everybody was surprised by the size and the abundance of the solder balls: 1-2mm (compared to the gap of 4.5mm between the pads) under more than 50% of the caps. Based on this observation, a decision was made on Monday to re-solder all the capacitors on the newly built wire planes. The above statistics applies for the two wire frames that were re-soldered by Anatoly so far. The leakage current after re-soldering went down to several nA for the whole frame. Not all the boards are like these: when we started testing the capacitors a month ago, Anatoly re-soldered the capacitors on another board, but we didn't see such solder balls there. After some discussions we agreed that it is worth informing the company. Chris will take care of the communications with them. Eugene: probably it's not worth pursuing this problem legally, but suggested to consult with the procurement.

- We will spend some time tomorrow preparing for the hurricane. The main problem we expect is the lack of electricity. Therefore, no gluing will be done tomorrow. Bill suggested to cover all the frames that are outside with vacuum bagging material.


- Bill wrote the procedure for the package assembly (linked), so far for two cells, and will continue working on it. Three types of cooling brackets will be attached to the pre-amps and that's why he suggested marking the corresponding daughter cards with different colors.

- Bill showed pictures (linked) made by Olga from the sample done after the last wire deadening. In the transition region the axises of the thin and thick part of the wire do not coincide which is strange, but other than that the diameters and the surface smoothness are OK.

- Gusset ring cut: Eugene explained that the main problem with the high currents (estimated as 7A in case of normal magnet quenching) on the gusset rings is the force that can reach several pounds directed mostly radially, but especially for the last package may have also z-component. Even if these forces are not so big it is worth cutting the rings because: the quench could be faster (it could damage the magnet as well but we can save the FDC at least); also these forces may change the positioning of the FDC which will require new calibrations.


- Chris finished testing the daughter boards (~500 in total), found 7 with problems and fixed most of them but two most likely will be unusable.

Chamber testing at 126 (cell #2)

- Beni showed results (page 573 from the FDC logbook) with the new gas mixture (pre-mixed bottle 89.8/10.2 Ar/CO2) and now the drift time spectrum looks much flatter than before (when using 99.9/11.1 ccpm mixture from two bottles). Still the spectrum is not as flat as it was with the prototype, most likely because the HV is not high enough.

- Beni made measurements (page 574, 575) with the same HV (+1975V) applied on both field and sense wires; this is to study the HV sector 4 on which we would not apply standard HV configuration. On some of the wires the efficiency is lower but it is unlikely that we have broken wires there.


A movie is linked above, that shows how fast we should produce a package.